### v1

# Cavity-Altered Superconductivity

Cavity-Altered Superconductivity
Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment? Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [9, 10, 11, 12]. Here, we devised and implemented a novel platform to realize cavity-altered materials. Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and superior quality factor [13, 14, 15, 16, 17, 18, 19]. We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET). The frequencies of infrared (IR) hyperbolic modes of hBN [20, 21] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [22, 23]. Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET. Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface. Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the superfluid suppression. These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state. This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground-state properties of materials using IR-active phonons.
Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice. The emergent effects born out of such strong interactions are abundant and drive the formation of a plethora of varied electronic and magnetic phases, as well as competing orders. At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter. A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring its quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8]. Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase of the electron phonon coupling via phonon polaritons [1, 18, 7], Amperean pairing directly mediated through dark-cavity photons [24, 25], and the driven-cavity extension of the Eliashberg effect [26]. Experiments with dark cavities have revealed altered ground states in the integer and fractional quantum Hall states of a 2D electron gas [9, 27, 28], as well as in charge density wave [10] systems, with additional reports suggesting changes to the diamagnetic response in superconducting nanoparticles [12, 11].
We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings. We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts with transition temperature () near 11.5 K [29, 30] (Supplementary Information Sections 1,2). We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN), interfaced with -ET (Fig. 1 bottom). Mode coupling at the interface is supported by the divergent photonic density of states [31, 32, 33] in the frequency range of hBN’s hyperbolicity (the permittivity has opposite signs along different axes [20, 21]), spanning the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [34]. Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode that is implicated in superconductivity [22, 23]. In this work, we examine how superconductivity is altered at the hBN/-ET interface, using scanning probe-based techniques to investigate local properties that are both inaccessible to bulk probes, and many surface probes due to the buried interfacial system.
Meissner force probe of resonant and non-resonant interfaces
Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [36, 37, 38, 39]. The force results from the supercurrent screening the stray magnetic field of the tip, making MFM sensitive to the superfluid density near the superconductor surface, where “near” signifies the decay length for currents running parallel to the surface . Notably, the Meissner force is not affected by inserting nonmagnetic materials between the tip and the superconductor, and data in Fig. 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers. We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40, 41], and analysis of allows extraction of the superfluid density [38] (Supplementary Information Section 4).
We examine the impact of hBN on superconductivity in a heterostructure composed of a 60 nm thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick). We report measured as a function of , the separation between the tip and the -ET surface, in Fig. 1a. On a region of the -ET crystal un-obscured by hBN (green), at T=2 K, is negative, indicating a repulsive force. The magnitude of grows with reduced tip-sample surface separation. The measurement above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading. As a control experiment, we placed a 55 nm thick RuCl3 microcrystal on top of -ET. RuCl3 is an insulator with a static permittivity similar to hBN [42, 35]. However all optical phonons of RuCl3 occur at much lower frequencies below 350 cm-1 [35] and therefore are not expected to resonantly couple to the C=C stretching mode of -ET (Fig. 1c schematically depicts the permittivities of hBN, -ET, and RuCl3). Comparing measurements above RuCl3/-ET (Fig. 1b, magenta curve) and bare -ET (green curve - same as in a), we observe a much weaker (less than %) effect than over hBN/-ET. The stark contrast between the RuCl3/-ET and hBN/-ET interfaces suggests that the suppression of superfluid density under hBN is a result of frequency matching across the hBN/-ET interface. While hBN and RuCl3 may exert mechanical forces on the -ET surface, superfluid density has been shown to be minimally impacted under pressure [43]. We perform a second control measurement on a heterostructure of 8 nm hBN on a 28.5 nm thick microcrystal of (BSCCO) with K hosting phonons below 650 cm-1 [44], far below the hyperbolic modes in hBN (Supplementary Information Section 5). In Fig. 1b (inset), hBN/BSCCO (blue) is indistinguishable from bare BSCCO (dark gray). Thus, the significant suppression of the superfluid density is unique to the hBN/-ET interface.
Electrodynamics of the hBN/-ET interface
Next, we investigate resonant mode coupling across the hBN/-ET interface. The hBN crystal hosts hyperbolic phonon polaritons (HPhPs) that exhibit light-like dispersion near the transverse optical (TO) frequency and phonon-like dispersion near the longitudinal optical (LO) frequency [13]. Here, TO and LO refer specifically to the in-plane modes. We examine the HPhP dispersion upon interfacing with -ET through the calculated imaginary part of the p-polarized Fresnel reflection , as shown in Fig. 2a (Supplementary Information Section 6). For free-standing hBN, the HPhP dispersion has a series of branches between the TO and LO phonon frequencies, each with a continuous dispersion [20]. By contrast, for an hBN/-ET heterostructure, the HPhP branches are interrupted near the C=C stretching frequency. Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully opened due to broadening caused by dielectric loss associated with the in-plane metallic response of -ET [46, 47].
To experimentally examine the mode coupling at the hBN/-ET interface, we use scattering-type scanning near-field optical microscopy (s-SNOM) [48] to evaluate the hBN HPhP dispersion. We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser. The strong electric field with high momentum achieved near the tip apex launches polariton waves. We performed nano-IR measurements at K on a region of a hBN/-ET heterostructure with 32 nm hBN on 100 nm -ET. Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images, as shown in Fig. 2b for cm-1. Profiles through s-SNOM images collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig. 2c). In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with frequency [49]. However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the dashed line tracking the fifth peak in Fig 2c.
To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode. First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP evanescent electric field with the C=C stretching mode. The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig 2d and Supplementary Information Section 2). Therefore, the C=C stretching mode does not effectively couple to the in-plane moments of the bare LO and TO phonons in hBN. In contrast, the HPhP evanescent field has a large out-of-plane component (Fig. 2d). To study the effect of this evanescent field, and its role in the renormalization of the zero-point motion of the C=C stretching mode, we calculated the C=C amplitude in the presence of an oscillating out-of-plane electric field induced by a phonon-polariton mode with a frequency spectrum (fundamental frequency cm-1), spanning the C=C frequency. With the temperature fixed at K, Langevin dynamics (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude, seen as a reduced amplitude and splitting of the peak in Fig. 2e. The negligible contribution of the direct phonon-phonon interaction is demonstrated by suppressing phonon-phonon scattering. Altogether, our Langevin molecular dynamics simulations indicate that cavity electrodynamics are crucial to explain the coupling between hBN and -ET. Importantly, vacuum fluctuations (zero-point phonon motion) and virtual processes invoked in mode hybridization permit the excitation of HPhPs without external photons, enabling coupling in the dark.
Suppression of superfluid density at the hBN/-ET interface
We confirm the association of reduced (Fig 1a) with suppressed superfluid density by studying the temperature dependence. Over bare -ET, the large repulsive signal is well described by the Meissner effect (Supplementary Information Section 4), decreases with increasing temperature (Fig 3a) and vanishes near . Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effect at the heterostructure interfaces by performing constant height scans with the tip 300 nm above the -ET surface in a 1616 m2 region including both hBN and RuCl3 (Fig. 3b-d). To facilitate comparison across temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET. At 2 K (panel b), pN/m over hBN indicates a weaker Meissner force. Similar to the curves in Fig. 1(a,b), the effect over RuCl3 is negligible. Warming to 12 K, just above , the image contrast disappears, giving everywhere. The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density. Further evaluation of measurement artifacts is available in Supplementary Information Section 8.
We have studied multiple hBN/-ET devices that show a suppression of superfluid density under hBN. The effect varies spatially (Supplementary Information Section 9); we focus on regions with the strongest suppression. The results are summarized in Fig. 4. To model the Meissner force analytically [36], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth (schematic in Fig. 4 inset) and evaluate the effective superfluid density under hBN, . In Fig. 4, we compile the suppression observed across all devices as a function of the local hBN (or RuCl3) thickness, where is the superfluid density in bare -ET. The minimum value of extracted on each device (maximal suppression) is at least 50%, emphasizing the sizeable effect. This result holds for a large variation of the hBN thicknesses ranging from 25 nm to 110 nm. Accurate conversion between and superfluid density would require knowledge of the depth profile. While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate between possible depth profiles all leading to similar Meissner force. However, the magnitude of the observed Meissner force suppression implies that a significant thickness of -ET is impacted, i.e. a depth at least the same order of magnitude as . A lower bound on the depth of suppression is roughly 430 nm (Supplementary Information Section 10). In conclusion, our hBN/-ET structures reveal the suppression of superfluid density in the presence of resonant mode coupling across the interface.
Cavity electrodynamics
Our findings indicate that cavity electrodynamics plays a role in the suppression of superfluid density. Specifically: (1) The C=C stretching mode intersects the hBN HPhP dispersion in a regime of highly mixed light-like and phonon-like character. (2) Significant coupling to the C=C stretching mode dipole requires the out-of-plane evanescent field of the HPhP, which typically extends several hundred nanometers outside hBN. The poor screening of interlayer electric fields in -ET [34] therefore opens the possibility for coupling deep within the crystal. Consequently, it is plausible that dark cavity electrodynamics impact the superconducting ground state through coupling to the C=C stretching mode.
Our simulations indicate that hBN cavity modes reduce C=C stretching mode fluctuations, suggesting that enhancing fluctuations, e.g. as previously suggested through mode softening in the cavity [1], may increase in -ET. The feasibility of a cavity-coupled ground state in -ET highlights the potential of dark cavities in engineering quantum electronic ground states through the control of the quantum dynamics in the material via coupling to engineered cavity vacuum fluctuations [19, 50, 16, 18].
Although our results apply to the -ET superconducting ground state, the concept of employing tunable vdW hyperbolic cavities to resonantly couple materials to cavity modes is highly general. vdW hyperbolic cavities host modes across the IR and THz energy ranges [15], allowing resonant coupling to excitations across this range, and in turn, ground state engineering of many electronic orders.
- [1] Sentef, M. A., Ruggenthaler, M. & Rubio, A. Cavity quantum-electrodynamical polaritonically enhanced electron-phonon coupling and its influence on superconductivity. Sci. Adv. 4, eaau6969 (2018).
- [2] Ruggenthaler, M., Tancogne-Dejean, N., Flick, J., Appel, H. & Rubio, A. From a quantum-electrodynamical light–matter description to novel spectroscopies. Nat. Rev. Chem. 2, 1–16 (2018).
- [3] Garcia-Vidal, F. J., Ciuti, C. & Ebbesen, T. W. Manipulating matter by strong coupling to vacuum fields. Science 373, eabd0336 (2021).
- [4] Latini, S. et al. The ferroelectric photo ground state of srtio3: Cavity materials engineering. PNAS 118, e2105618118 (2021).
- [5] Schlawin, F., Kennes, D. M. & Sentef, M. A. Cavity quantum materials. Appl. Phys. Rev. 9 (2022).
- [6] Viñas Boström, E., Sriram, A., Claassen, M. & Rubio, A. Controlling the magnetic state of the proximate quantum spin liquid -RuCl3 with an optical cavity. npj Comput. Mater. 9, 202 (2023).
- [7] Lu, I.-T. et al. Cavity-enhanced superconductivity in MgB2 from first-principles quantum electrodynamics (QEDFT). PNAS 121, e2415061121 (2024).
- [8] Lu, I.-T. et al. Cavity engineering of solid-state materials without external driving. Adv. Opt. Photon. 17, 441–525 (2025).
- [9] Paravicini-Bagliani, G. et al. Magneto-transport controlled by landau polariton states. Nat. Phys. 15, 186–190 (2019).
- [10] Jarc, G. et al. Cavity-mediated thermal control of metal-to-insulator transition in 1T-TaS2. Nature 622, 487–492 (2023).
- [11] Thomas, A. et al. Large Enhancement of Ferromagnetism under a Collective Strong Coupling of YBCO Nanoparticles. Nano Lett. 21, 4365–4370 (2021).
- [12] Thomas, A. et al. Exploring superconductivity under strong coupling with the vacuum electromagnetic field. J. Chem. Phys. 162, 134701 (2025).
- [13] Basov, D., Fogler, M. & García de Abajo, F. Polaritons in van der waals materials. Science 354, aag1992 (2016).
- [14] Low, T. et al. Polaritons in layered two-dimensional materials. Nat. Mater. 16, 182–194 (2017).
- [15] Basov, D. N., Asenjo-Garcia, A., Schuck, P. J., Zhu, X. & Rubio, A. Polariton panorama. Nanophotonics 10, 549–577 (2020).
- [16] Bloch, J., Cavalleri, A., Galitski, V., Hafezi, M. & Rubio, A. Strongly correlated electron–photon systems. Nature 606, 41–48 (2022).
- [17] Ebbesen, T. W., Rubio, A. & Scholes, G. D. Introduction: Polaritonic Chemistry. Chem. Rev. 123, 12037–12038 (2023).
- [18] Hübener, H., Boström, E. V., Claassen, M., Latini, S. & Rubio, A. Quantum materials engineering by structured cavity vacuum fluctuations. Mater. Quantum. Technol. 4, 023002 (2024).
- [19] Herzig Sheinfux, H. et al. High-quality nanocavities through multimodal confinement of hyperbolic polaritons in hexagonal boron nitride. Nat. Mater. 23, 499–505 (2024).
- [20] Dai, S. et al. Tunable phonon polaritons in atomically thin van der waals crystals of boron nitride. Science 343, 1125–1129 (2014).
- [21] Caldwell, J. D. et al. Sub-diffractional volume-confined polaritons in the natural hyperbolic material hexagonal boron nitride. Nat. Commun. 5, 5221 (2014).
- [22] Buzzi, M. et al. Photomolecular High-Temperature Superconductivity. Phys. Rev. X. 10, 31028 (2020).
- [23] Buzzi, M. et al. Phase Diagram for Light-Induced Superconductivity in . Phys. Rev. Lett. 127, 197002 (2021).
- [24] Schlawin, F., Cavalleri, A. & Jaksch, D. Cavity-mediated electron-photon superconductivity. Phys. Rev. Lett. 122, 133602 (2019).
- [25] Andolina, G. M. et al. Amperean superconductivity cannot be induced by deep subwavelength cavities in a two-dimensional material. Phys. Rev. B 109, 104513 (2024).
- [26] Curtis, J. B., Raines, Z. M., Allocca, A. A., Hafezi, M. & Galitski, V. M. Cavity quantum eliashberg enhancement of superconductivity. Phys. Rev. Lett. 122, 167002 (2019).
- [27] Enkner, J. et al. Testing the renormalization of the von klitzing constant by cavity vacuum fields. Phys. Rev. X 14, 021038 (2024).
- [28] Enkner, J. et al. Tunable vacuum-field control of fractional and integer quantum hall phases. Nature 641, 884–889 (2025).
- [29] Ishiguro, T., Yamaji, K. & Saito, G. Organic superconductors (Springer-Verlag, Berlin, Heidelberg, 1998), 2nd edn.
- [30] Ardavan, A. et al. Recent Topics of Organic Superconductors. J. Phys. Soc. Jpn. 81, 11004 (2012).
- [31] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photonics 7, 948–957 (2013).
- [32] Sun, J., Litchinitser, N. M. & Zhou, J. Indefinite by Nature: From Ultraviolet to Terahertz. ACS Photonics 1, 293–303 (2014).
- [33] Kim, J. et al. Improving the radiative decay rate for dye molecules with hyperbolic metamaterials. Opt. Express 20, 8100–8116 (2012).
- [34] McGuire, J. J. et al. Incoherent interplane conductivity of . Phys. Rev. B 64, 094503 (2001).
- [35] Reschke, S. et al. Electronic and phonon excitations in -RuCl3. Phys. Rev. B 96, 165120 (2017).
- [36] Xu, J., Miller Jr, J. & Ting, C. Magnetic levitation force and penetration depth in type-II superconductors. Phys. Rev. B 51, 424 (1995).
- [37] Roseman, M., Grütter, P., Badia, A. & Metlushko, V. Flux lattice imaging of a patterned niobium thin film. J. Appl. Phys. 89, 6787–6789 (2001).
- [38] Luan, L. et al. Local measurement of the penetration depth in the pnictide superconductor . Phys. Rev. B 81, 100501 (2010).
- [39] Lamhot, Y. et al. Local characterization of superconductivity in . Phys. Rev. B 91, 060504 (2015).
- [40] Schwarz, A. & Wiesendanger, R. Magnetic sensitive force microscopy. Nano Today 3, 28–39 (2008).
- [41] Hartmann, U. Magnetic Force Microscopy. Annu. Rev. Mater. Sci. 29, 53–87 (1999).
- [42] Laturia, A., Van de Put, M. L. & Vandenberghe, W. G. Dielectric properties of hexagonal boron nitride and transition metal dichalcogenides: from monolayer to bulk. npj 2D Mater. Appl. 2, 6 (2018).
- [43] Wakamatsu, K., Miyagawa, K. & Kanoda, K. Superfluid density versus transition temperature in a layered organic superconductor -(BEDT-TTF)2Cu[N(CN)2]Br under pressure. Phys. Rev. Res. 2, 043008 (2020).
- [44] Kovaleva, N. et al. c-axis lattice dynamics in Bi-based cuprate superconductors. Phys. Rev. B 69, 054511 (2004).
- [45] Vicentini, E. et al. Pseudoheterodyne interferometry for multicolor near-field imaging. Opt. express 31, 22308–22322 (2023).
- [46] Sedlmeier, K. et al. Absence of charge order in the dimerized -phase BEDT-TTF salts. Phys. Rev. B 86, 245103 (2012).
- [47] Faltermeier, D. et al. Bandwidth-controlled mott transition in : Optical studies of localized charge excitations. Phys. Rev. B 76, 165113 (2007).
- [48] Chen, X. et al. Modern scattering-type scanning near-field optical microscopy for advanced material research. Adv. Mater. 31, 1804774 (2019).
- [49] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photonics 15, 197–202 (2021).
- [50] Le Dé, B., Eckhardt, C. J., Kennes, D. M. & Sentef, M. A. Cavity engineering of hubbard u via phonon polaritons. J.Phys. Mater. 5, 024006 (2022).
Methods
Device fabrication
Heterostructures were fabricated by placing vdW materials, on the order of 10 nm to 120 nm thick, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2). Measurements were performed on five devices (Supplementary Information Section 3). Data in Fig 1 is from device 1, Fig 2 from device 2, Fig 3(b-d) from device 1, and Fig 3a from device 5. Fig 4 is a compilation of results from devices 1, 3, and 4.
The -ET crystal surfaces were either as-grown (devices 1 and 3), exposed by cleaving bulk crystals (devices 4 and 5), or exposed by exfoliating single crystals (device 2). For bulk samples, the -ET single crystal was fixed onto a silicon substrate using either epoxy or silver paint. For devices 4 and 5, the crystal was then cleaved to expose new surfaces. Large atomically flat surfaces were then selected by optical and atomic force microscopy. As-grown surfaces were regularly found to be atomically flat. Cleaved surfaces showed larger variations in surface roughness.
Both hBN and RuCl3 flakes were exfoliated using a standard scotch-tape technique onto a polydimethylsiloxane (PDMS) stamp and identified via optical and atomic force microscopy. The stacking process was performed using a commercial transfer stage with temperature control. The PDMS stamp was gently pressed onto the -ET surface and then slowly retracted. The hBN flakes were released at approximately 60°C, while the RuCl3 flakes were released at around 110°C. Isotopically pure hBN (10BN) was used to achieve a narrow linewidth for the TO phonon. The narrow linewidth makes HPhPs visible in s-SNOM.
hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN ctystals in a golvebox filled with inert gas ( ppm; ppm). BSCCO was first exfoliated on SiO2 substrates. hBN flakes were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
s-SNOM
The nano-infrared scattering experiments were performed using a home-built cryogenic scattering-type scanning near-field optical microscope (s-SNOM) housed in an ultra-high vacuum chamber with a base pressure of torr. The s-SNOM is based on a tapping-mode atomic force microscope (AFM) with the tapping frequency and amplitude of the AFM set to approximately 285 kHz and 70 nm, respectively. The s-SNOM operates by scattering tightly focused laser light from a sharp AFM tip, enabling near-field interactions to be probed with nanometer-scale resolution. The laser source is a tunable quantum cascade laser from Daylight Solutions. The laser beam was focused onto the metallized AFM tip using a parabolic mirror with a 12-mm focal length. The back-scattered light was registered by a mercury cadmium telluride detector and demodulated following a pseudoheterodyne scheme. The signal was demodulated at the th harmonic of the tip tapping frequency, yielding background-free near-field signal. In this work, we selected =3 and 4 to eliminate the far-field background.
MFM
MFM measurements were performed in an Attocube cantilever-based cryogenic AFM (attoAFM I with attoLIQUID 2000 cryostat), where the microscope sits in a helium exchange gas at low temperature. Nanosensors PPP-MFMR probes with hard magnetic coating and force constants near N/m were used for all measurements. Measurements were taken on different samples using different tips, all with resonance frequencies ranging from 75.5 kHz to 79.4 kHz. Measurements recorded the resonant frequency shift of the probe cantilever () and were converted to via [41], using the force constants provided by the manufacturer for each cantilever and the measured resonant frequency far from the sample surface, . The tip oscillation amplitude was typically 50 nm - 70 nm. hBN is non-magnetic and therefore does not impact the detection of the Meissner effect over hBN/-ET. RuCl3 is antiferromagnetic at low temperature, but has a low susceptibility.
To minimize noise and instrument drift, was typically averaged over multiple measurements repeated at each point and taken with high resolution (typically 1-2 nm to define the surface accurately). During the measurements, the tip was electrically biased to compensate for the local contact potential difference and thus minimize electrostatic forces. The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise. For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the Meissner force is negligible. For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 5). For proper comparison in Fig 1, we have compensated for the thickness of the hBN and RuCl3 such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
Constant height MFM images were taken with the tip at a fixed height above the plane of the -ET surface. The images shown are raw data, except for the linear conversion from to and an overall constant offset to calculate . During the measurements, the tip was held at a fixed electric bias to compensate for the contact potential difference at one point in the field of view.
The data points in Fig. 4 were obtained either by fitting curves or by converting constant height MFM images to maps of the local , as described in Supplementary Information Section 4.
HPhP dispersion simulation
The HPhP dispersion was calculated using the transfer matrix method. The dispersion is represented by peaks in the imaginary part of the Fresnel reflectivity of the device for p-polarized light, . In the dispersion calculation, we consider the real-valued frequencies and momentum. For coupling analysis, the eigenmodes are found by determining the poles of , where the complex frequencies and real momentum are considered.

### v3

@@
 Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment?
-Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [9, 10, 11, 12].
+Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [10, 11, 12, 13, 14, 15].
 Here, we devised and implemented a novel platform to realize cavity-altered materials.
-Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and superior quality factor [13, 14, 15, 16, 17, 18, 19].
+Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [16, 17, 18, 19, 20, 21, 22].
 We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET).
 The frequencies of infrared (IR) hyperbolic modes of hBN [23, 24] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [25, 26].
-Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET.
+Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm the presence of resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET.
 Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface.
@@
 These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state.
-This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground-state properties of materials using IR-active phonons.
+This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties.
 Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice.
-The emergent effects born out of such strong interactions are abundant and drive the formation of a plethora of varied electronic and magnetic phases, as well as competing orders.
+The emergent effects born out of such strong interactions are abundant and drive the formation of varied electronic and magnetic phases.
 At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter.
-A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring its quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8].
-Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase of the electron phonon coupling via phonon polaritons [1, 18, 7], Amperean pairing directly mediated through dark-cavity photons [24, 25], and the driven-cavity extension of the Eliashberg effect [26].
-Experiments with dark cavities have revealed altered ground states in the integer and fractional quantum Hall states of a 2D electron gas [9, 27, 28], as well as in charge density wave [10] systems, with additional reports suggesting changes to the diamagnetic response in superconducting nanoparticles [12, 11].
+A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9].
+Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 21, 7]; Amperean pairing directly mediated through dark-cavity photons [27, 28]; and the driven-cavity extension of the Eliashberg effect [29].
+Experiments with dark cavities revealed modified transport properties in the integer and fractional quantum Hall states of a 2D electron gas [10, 13, 15] as well as cavity-assisted thermal control of the metal-to-insulator transition in charge density wave systems [11].
 We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings.
-We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts with transition temperature () near 11.5 K [29, 30] (Supplementary Information Sections 1,2).
-We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN), interfaced with -ET (Fig.
-1 bottom).
-Mode coupling at the interface is supported by the divergent photonic density of states [31, 32, 33] in the frequency range of hBN’s hyperbolicity (the permittivity has opposite signs along different axes [20, 21]), spanning the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [34].
-Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode that is implicated in superconductivity [22, 23].
-In this work, we examine how superconductivity is altered at the hBN/-ET interface, using scanning probe-based techniques to investigate local properties that are both inaccessible to bulk probes, and many surface probes due to the buried interfacial system.
+We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts below the transition temperature =11.5 K [30, 31] (Supplementary Information Sections 1,2).
+We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN).
+Hyperbolicity occurs when the permittivity has opposite signs along different axes [23, 24] and results in a highly enhanced photonic density of states [32, 33, 34].
+Notably, hyperbolic modes of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [35] and the two modes hybridize at the hBN/-ET interface.
+Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [25, 26].
+In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques that effectively analyze the electrodynamics of buried interfaces.
 Meissner force probe of resonant and non-resonant interfaces
-Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [36, 37, 38, 39].
-The force results from the supercurrent screening the stray magnetic field of the tip, making MFM sensitive to the superfluid density near the superconductor surface, where “near” signifies the decay length for currents running parallel to the surface .
-Notably, the Meissner force is not affected by inserting nonmagnetic materials between the tip and the superconductor, and data in Fig.
+Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [37, 38, 39].
+The force results from the supercurrent screening the stray magnetic field of the tip.
+Notably, the Meissner force is not affected when non-magnetic materials are inserted between the tip and the superconductor, and the data in Fig.
 1(a,b) attest that MFM probes superconductors encapsulated by insulating layers.
-We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40, 41], and analysis of allows extraction of the superfluid density [38] (Supplementary Information Section 4).
-We examine the impact of hBN on superconductivity in a heterostructure composed of a 60 nm thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick).
+We measure the shift in the resonant frequency of the cantilever, which is proportional to the force gradient [40].
+Analysis of allows the extraction of the superfluid density [38] (Methods and Supplementary Information Section 4).
+We examine the impact of hBN on superconductivity in a representative heterostructure composed of a 60-nm-thick hBN microcrystal on the surface of a bulk -ET single crystal (200 thick).
 We report measured as a function of , the separation between the tip and the -ET surface, in Fig.
 1a.
-On a region of the -ET crystal un-obscured by hBN (green), at T=2 K, is negative, indicating a repulsive force.
+On a region of the -ET crystal unobscured by hBN (green), is negative at K, indicating a repulsive force.
 The magnitude of grows with reduced tip-sample surface separation.
-The measurement above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading.
-As a control experiment, we placed a 55 nm thick RuCl3 microcrystal on top of -ET.
-RuCl3 is an insulator with a static permittivity similar to hBN [42, 35].
-However all optical phonons of RuCl3 occur at much lower frequencies below 350 cm-1 [35] and therefore are not expected to resonantly couple to the C=C stretching mode of -ET (Fig.
-1c schematically depicts the permittivities of hBN, -ET, and RuCl3).
+The signal collected above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading.
+As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of -ET.
+RuCl3 is an insulator with a static permittivity similar to that of hBN [41, 36].
+However, optical phonons of RuCl3 all occur at much lower frequencies below 350 cm-1 [36] and therefore do not resonantly couple to the C=C stretching mode of -ET.
+Fig.
+1c schematically depicts the permittivities of hBN, -ET, and RuCl3.
 Comparing measurements above RuCl3/-ET (Fig.
-1b, magenta curve) and bare -ET (green curve - same as in a), we observe a much weaker (less than %) effect than over hBN/-ET.
-The stark contrast between the RuCl3/-ET and hBN/-ET interfaces suggests that the suppression of superfluid density under hBN is a result of frequency matching across the hBN/-ET interface.
-While hBN and RuCl3 may exert mechanical forces on the -ET surface, superfluid density has been shown to be minimally impacted under pressure [43].
-We perform a second control measurement on a heterostructure of 8 nm hBN on a 28.5 nm thick microcrystal of (BSCCO) with K hosting phonons below 650 cm-1 [44], far below the hyperbolic modes in hBN (Supplementary Information Section 5).
-In Fig.
-1b (inset), hBN/BSCCO (blue) is indistinguishable from bare BSCCO (dark gray).
-Thus, the significant suppression of the superfluid density is unique to the hBN/-ET interface.
+1b, magenta curve) and bare -ET (green curve - same as in Fig.
+1a), we observe a much weaker (less than %) effect than over hBN/-ET.
+The stark contrast between the RuCl3/-ET and hBN/-ET interfaces indicates that the suppression of superfluid density under hBN is likely prompted by resonant coupling across the hBN/-ET interface.
+We remark that the RuCl3 control experiment rules out a prominent role for work function disparity [42] in the suppression of the superfluid density in Fig.
+1 and similarly suggests that strain is unlikely to be significant.
+We perform a second control measurement on a heterostructure of 8-nm hBN residing on a 28.5-nm-thick microcrystal of (BSCCO) with K.
+Crystals of BSCCO host phonons below 650 cm-1 [43], far below the hyperbolic modes in hBN (Supplementary Information Section 5).
+MFM data plotted in Fig.
+1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray).
+Thus, the significant suppression of the superfluid density is specific to the resonant nature of the hBN/-ET interface.
 Electrodynamics of the hBN/-ET interface
@@
 Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully opened due to broadening caused by dielectric loss associated with the in-plane metallic response of -ET [45, 46].
-To experimentally examine the mode coupling at the hBN/-ET interface, we use scattering-type scanning near-field optical microscopy (s-SNOM) [48] to evaluate the hBN HPhP dispersion.
+To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [47].
+Specifically, we evaluated the hBN HPhP dispersion as these modes couple to the C=C stretching mode of -ET.
 We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser.
 The strong electric field with high momentum achieved near the tip apex launches polariton waves.
-We performed nano-IR measurements at K on a region of a hBN/-ET heterostructure with 32 nm hBN on 100 nm -ET.
-Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images, as shown in Fig.
+We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET.
+Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig.
 2b for cm-1.
-Profiles through s-SNOM images collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig.
+Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig.
 2c).
-In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with frequency [49].
-However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the dashed line tracking the fifth peak in Fig 2c.
+In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [48].
+However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig.
+2c.
+These nano-IR data attest to mode coupling across the hBN--ET interface.
 To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode.
-First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP evanescent electric field with the C=C stretching mode.
-The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig 2d and Supplementary Information Section 2).
-Therefore, the C=C stretching mode does not effectively couple to the in-plane moments of the bare LO and TO phonons in hBN.
-In contrast, the HPhP evanescent field has a large out-of-plane component (Fig.
+First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP’s electric field with the C=C stretching mode.
+The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig.
+2d and Supplementary Information Section 2).
+It is important to note that the HPhP field has an out-of-plane component that can couple to the C=C mode (Fig.
 2d).
-To study the effect of this evanescent field, and its role in the renormalization of the zero-point motion of the C=C stretching mode, we calculated the C=C amplitude in the presence of an oscillating out-of-plane electric field induced by a phonon-polariton mode with a frequency spectrum (fundamental frequency cm-1), spanning the C=C frequency.
-With the temperature fixed at K, Langevin dynamics (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude, seen as a reduced amplitude and splitting of the peak in Fig.
+The combined contributions of many HPhP branches outspreading to high momenta make this field extend far beyond the hBN surface.
+We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1).
+Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude and splits the peak in Fig.
 2e.
-The negligible contribution of the direct phonon-phonon interaction is demonstrated by suppressing phonon-phonon scattering.
-Altogether, our Langevin molecular dynamics simulations indicate that cavity electrodynamics are crucial to explain the coupling between hBN and -ET.
-Importantly, vacuum fluctuations (zero-point phonon motion) and virtual processes invoked in mode hybridization permit the excitation of HPhPs without external photons, enabling coupling in the dark.
+We therefore conclude that vacuum fluctuations—specifically, zero-point phonon motion—mediate the coupling between hyperbolic phonon polaritons and molecular vibrations without the need for external photons, thus enabling interaction in the dark.
 Suppression of superfluid density at the hBN/-ET interface
-We confirm the association of reduced (Fig 1a) with suppressed superfluid density by studying the temperature dependence.
-Over bare -ET, the large repulsive signal is well described by the Meissner effect (Supplementary Information Section 4), decreases with increasing temperature (Fig 3a) and vanishes near .
-Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effect at the heterostructure interfaces by performing constant height scans with the tip 300 nm above the -ET surface in a 1616 m2 region including both hBN and RuCl3 (Fig.
+We now validate the association of reduced (Fig.
+1a) with suppressed superfluid density by studying the temperature dependence.
+Over bare -ET, the large repulsive is rooted in the Meissner effect (Supplementary Information Section 4), it decreases with increasing temperature (Fig.
+3a) and vanishes near .
+Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces by performing constant-height scans with the tip hovering above the -ET surface in a 1616-m2 region that includes both hBN and RuCl3 (Fig.
 3b-d).
-To facilitate comparison across temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET.
-At 2 K (panel b), pN/m over hBN indicates a weaker Meissner force.
-Similar to the curves in Fig.
+To facilitate comparison between temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET.
+At 2 K (panel b), pN/m over hBN, indicating a weaker Meissner force.
+As with the curves in Fig.
 1(a,b), the effect over RuCl3 is negligible.
@@
 The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density.
-Further evaluation of measurement artifacts is available in Supplementary Information Section 8.
-We have studied multiple hBN/-ET devices that show a suppression of superfluid density under hBN.
-The effect varies spatially (Supplementary Information Section 9); we focus on regions with the strongest suppression.
+Further evaluation and additional control measurements are described in Supplementary Information Section 8.
+We observe suppression of superfluid density under hBN in multiple hBN/-ET heterostructures.
+The magnitude of suppression varies spatially (Supplementary Information Section 9); we focus on regions with the strongest effect.
 The results are summarized in Fig.
 4.
-To model the Meissner force analytically [36], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth (schematic in Fig.
-4 inset) and evaluate the effective superfluid density under hBN, .
+To model the Meissner force analytically [37], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, .
 In Fig.
-4, we compile the suppression observed across all devices as a function of the local hBN (or RuCl3) thickness, where is the superfluid density in bare -ET.
-The minimum value of extracted on each device (maximal suppression) is at least 50%, emphasizing the sizeable effect.
-This result holds for a large variation of the hBN thicknesses ranging from 25 nm to 110 nm.
-Accurate conversion between and superfluid density would require knowledge of the depth profile.
-While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate between possible depth profiles all leading to similar Meissner force.
-However, the magnitude of the observed Meissner force suppression implies that a significant thickness of -ET is impacted, i.e.
-a depth at least the same order of magnitude as .
-A lower bound on the depth of suppression is roughly 430 nm (Supplementary Information Section 10).
-In conclusion, our hBN/-ET structures reveal the suppression of superfluid density in the presence of resonant mode coupling across the interface.
-Cavity electrodynamics
-Our findings indicate that cavity electrodynamics plays a role in the suppression of superfluid density.
-Specifically: (1) The C=C stretching mode intersects the hBN HPhP dispersion in a regime of highly mixed light-like and phonon-like character.
-(2) Significant coupling to the C=C stretching mode dipole requires the out-of-plane evanescent field of the HPhP, which typically extends several hundred nanometers outside hBN.
-The poor screening of interlayer electric fields in -ET [34] therefore opens the possibility for coupling deep within the crystal.
-Consequently, it is plausible that dark cavity electrodynamics impact the superconducting ground state through coupling to the C=C stretching mode.
-Our simulations indicate that hBN cavity modes reduce C=C stretching mode fluctuations, suggesting that enhancing fluctuations, e.g.
-as previously suggested through mode softening in the cavity [1], may increase in -ET.
-The feasibility of a cavity-coupled ground state in -ET highlights the potential of dark cavities in engineering quantum electronic ground states through the control of the quantum dynamics in the material via coupling to engineered cavity vacuum fluctuations [19, 50, 16, 18].
-Although our results apply to the -ET superconducting ground state, the concept of employing tunable vdW hyperbolic cavities to resonantly couple materials to cavity modes is highly general.
-vdW hyperbolic cavities host modes across the IR and THz energy ranges [15], allowing resonant coupling to excitations across this range, and in turn, ground state engineering of many electronic orders.
+4, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses.
+We present data in the form of , where is the superfluid density in bare -ET.
+In all heterostructures, we observe at least 50% suppression of the superfluid density prompted by the hBN cavity.
+This result holds over a large variation in the hBN thickness, ranging from 25 nm to 110 nm.
+Accurate conversion between and superfluid density would require knowledge of the depth profile (Supplementary Information Section 4).
+While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar Meissner forces.
+However, the magnitude of the observed Meissner force suppression implies that -ET is impacted by the cavity to a significant depth.
+The impacted depth appears to be of the same order of magnitude as the polaritonic wavelength (Supplementary Information Section 10).
+Hyperbolic cavity electrodynamics
+The data in Figs.
+1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium property of superconductivity.
+These experiments establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering.
+Our specific implementation of the resonant structure reveals two notable features of hyperbolic cavity control.
+First, hyperbolic modes entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields.
+The out-of-plane component is crucial for coupling to the out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation.
+Second, the multiple branches of hyperbolic modes span a broad momentum range, enabling continuous interaction with the weakly dispersing C=C resonance across the entire phase space, as shown in Fig.
+2a.
+Future experiments will be needed to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [49].
+Our data and modeling suggest a plausible pathway for cavity enhancement of superconducting properties.
+Indeed, our simulations indicate that the hBN cavity subdues the C=C stretching mode fluctuations prompting the stark suppression of superfluid density.
+By the same token, boosting fluctuations through mode softening in the cavity [1] may increase the superfluid density and possibly in -ET.
+Furthermore, the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulations of material properties is rather general.
+We remind the reader that vdW crystals host hyperbolic modes of phononic, plasmonic and exitonic origin [50] spanning the range from THz to visible, thus offering limitless options for hyper-cavity engineering.
 - [1] Sentef, M.
@@
 - [7] Lu, I.-T. et al. Cavity-enhanced superconductivity in MgB2 from first-principles quantum electrodynamics (QEDFT). PNAS 121, e2415061121 (2024).
+- [8] Flick, J., Rivera, N. & Narang, P. Strong light-matter coupling in quantum chemistry and quantum photonics. Nanophotonics 7, 1479–1501 (2018).
 - [9] Lu, I.-T. et al. Cavity engineering of solid-state materials without external driving. Adv. Opt. Photon. 17, 441–525 (2025).
@@
 - [12] Thomas, A. et al. Large Enhancement of Ferromagnetism under a Collective Strong Coupling of YBCO Nanoparticles. Nano Lett. 21, 4365–4370 (2021).
+- [13] Enkner, J. et al. Testing the renormalization of the von klitzing constant by cavity vacuum fields. Phys. Rev. X 14, 021038 (2024).
 - [14] Thomas, A. et al. Exploring superconductivity under strong coupling with the vacuum electromagnetic field. J. Chem. Phys. 162, 134701 (2025).
+- [15] Enkner, J. et al. Tunable vacuum-field control of fractional and integer quantum hall phases. Nature 641, 884–889 (2025).
 - [16] Basov, D., Fogler, M. & García de Abajo, F. Polaritons in van der waals materials. Science 354, aag1992 (2016).
@@
 - [29] Curtis, J. B., Raines, Z. M., Allocca, A. A., Hafezi, M. & Galitski, V. M. Cavity quantum eliashberg enhancement of superconductivity. Phys. Rev. Lett. 122, 167002 (2019).
-- [27] Enkner, J. et al. Testing the renormalization of the von klitzing constant by cavity vacuum fields. Phys. Rev. X 14, 021038 (2024).
-- [28] Enkner, J. et al. Tunable vacuum-field control of fractional and integer quantum hall phases. Nature 641, 884–889 (2025).
-- [29] Ishiguro, T., Yamaji, K. & Saito, G. Organic superconductors (Springer-Verlag, Berlin, Heidelberg, 1998), 2nd edn.
-- [30] Ardavan, A. et al. Recent Topics of Organic Superconductors. J. Phys. Soc. Jpn. 81, 11004 (2012).
+- [30] Kanoda, K., Sakao, K., Takahashi, T., Komatsu, T. & Saito, G. NMR study of an organic superconductor, -(BEDT-TTF)2Cu [N(CN)2]Br. Physica C. 185, 2667–2668 (1991).
+- [31] Dressel, M. et al. Electrodynamics of the organic superconductorsinebreak -(BEDT-TTF)2Cu(NCS)2 and -(BEDT-TTF)2Cu[N(CN)2]Br. Phys. Rev. B 50, 13603 (1994).
 - [32] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photonics 7, 948–957 (2013).
@@
 - [37] Xu, J., Miller Jr, J. & Ting, C. Magnetic levitation force and penetration depth in type-II superconductors. Phys. Rev. B 51, 424 (1995).
-- [37] Roseman, M., Grütter, P., Badia, A. & Metlushko, V. Flux lattice imaging of a patterned niobium thin film. J. Appl. Phys. 89, 6787–6789 (2001).
 - [38] Luan, L. et al. Local measurement of the penetration depth in the pnictide superconductor . Phys. Rev. B 81, 100501 (2010).
 - [39] Lamhot, Y. et al. Local characterization of superconductivity in . Phys. Rev. B 91, 060504 (2015).
-- [40] Schwarz, A. & Wiesendanger, R. Magnetic sensitive force microscopy. Nano Today 3, 28–39 (2008).
 - [40] Hartmann, U. Magnetic Force Microscopy. Annu. Rev. Mater. Sci. 29, 53–87 (1999).
@@
 2, 6 (2018).
-- [43] Wakamatsu, K., Miyagawa, K.
-& Kanoda, K.
-Superfluid density versus transition temperature in a layered organic superconductor -(BEDT-TTF)2Cu[N(CN)2]Br under pressure.
-Phys.
-Rev.
-Res.
-2, 043008 (2020).
+- [42] Rizzo, D. J. et al. Charge-transfer plasmon polaritons at graphene/-RuCl3 interfaces. Nano lett. 20, 8438–8445 (2020).
 - [43] Kovaleva, N. et al. c-axis lattice dynamics in Bi-based cuprate superconductors. Phys. Rev. B 69, 054511 (2004).
@@
 - [48] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photonics 15, 197–202 (2021).
-- [50] Le Dé, B., Eckhardt, C. J., Kennes, D. M. & Sentef, M. A. Cavity engineering of hubbard u via phonon polaritons. J.Phys. Mater. 5, 024006 (2022).
+- [49] Sternbach, A. J. et al. Negative refraction in hyperbolic hetero-bicrystals. Science 379, 555–557 (2023).
+- [50] Basov, D. et al. Polaritonic quantum matter. Nanophotonics (2025).
 Methods
@@
 Measurements were performed on five devices (Supplementary Information Section 3).
-Data in Fig 1 is from device 1, Fig 2 from device 2, Fig 3(b-d) from device 1, and Fig 3a from device 5.
-Fig 4 is a compilation of results from devices 1, 3, and 4.
+Data in Fig.
+1 are from device 1, in Fig.
+2 from device 2, in Fig.
+3(b-d) from device 1, and in Fig.
+3a from device 5.
+Fig.
+4 is a compilation of results from devices 1, 3, and 4.
 The -ET crystal surfaces were either as-grown (devices 1 and 3), exposed by cleaving bulk crystals (devices 4 and 5), or exposed by exfoliating single crystals (device 2).
@@
 The narrow linewidth makes HPhPs visible in s-SNOM.
-hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN ctystals in a golvebox filled with inert gas ( ppm; ppm).
+hBN/BSCCO samples were prepared by exfoliating and assembling BSCCO and hBN crystals in a glovebox filled with inert gas ( ppm; ppm).
 BSCCO was first exfoliated on SiO2 substrates.
@@
 The laser beam was focused onto the metallized AFM tip using a parabolic mirror with a 12-mm focal length.
-The back-scattered light was registered by a mercury cadmium telluride detector and demodulated following a pseudoheterodyne scheme.
-The signal was demodulated at the th harmonic of the tip tapping frequency, yielding background-free near-field signal.
+The backscattered light was registered by a mercury cadmium telluride detector and demodulated following a pseudoheterodyne scheme.
+The signal was demodulated at the th harmonic of the tip tapping frequency, yielding a background-free near-field signal.
 In this work, we selected =3 and 4 to eliminate the far-field background.
@@
 MFM measurements were performed in an Attocube cantilever-based cryogenic AFM (attoAFM I with attoLIQUID 2000 cryostat), where the microscope sits in a helium exchange gas at low temperature.
-Nanosensors PPP-MFMR probes with hard magnetic coating and force constants near N/m were used for all measurements.
+Nanosensors PPP-MFMR probes with hard magnetic coatings and force constants near N/m were used for all measurements.
 Measurements were taken on different samples using different tips, all with resonance frequencies ranging from 75.5 kHz to 79.4 kHz.
-Measurements recorded the resonant frequency shift of the probe cantilever () and were converted to via [41], using the force constants provided by the manufacturer for each cantilever and the measured resonant frequency far from the sample surface, .
+Measurements recorded the resonant frequency shift of the probe cantilever () and were converted to via [40], using the force constant provided by the manufacturer for each cantilever and the measured resonant frequency far from the sample surface, .
 The tip oscillation amplitude was typically 50 nm - 70 nm.
-hBN is non-magnetic and therefore does not impact the detection of the Meissner effect over hBN/-ET.
+hBN is non-magnetic and therefore has no impact on the detection of the Meissner effect over hBN/-ET.
 RuCl3 is antiferromagnetic at low temperature, but has a low susceptibility.
@@
 For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 5).
-For proper comparison in Fig 1, we have compensated for the thickness of the hBN and RuCl3 such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
-Constant height MFM images were taken with the tip at a fixed height above the plane of the -ET surface.
+For proper comparison in Fig.
+1, we have compensated for the thickness of the hBN and RuCl3 microcrystals such that the measurements for hBN/-ET, bare -ET, and RuCl3/-ET all have the -ET surface at zero.
+Constant-height MFM images were taken with the tip at a fixed height above the plane of the -ET surface.
 The images shown are raw data, except for the linear conversion from to and an overall constant offset to calculate .
 During the measurements, the tip was held at a fixed electric bias to compensate for the contact potential difference at one point in the field of view.
-The data points in Fig. 4 were obtained either by fitting curves or by converting constant height MFM images to maps of the local , as described in Supplementary Information Section 4.
+The data points in Fig.
+4 were obtained either by fitting curves or by converting constant-height MFM images to maps of the local , the decay length for currents running parallel to the surface.
+was converted to superfluid density via the proportionality to , as described in Supplementary Information Section 4.
 HPhP dispersion simulation
@@
 For coupling analysis, the eigenmodes are found by determining the poles of , where the complex frequencies and real momentum are considered.
+Data availability
+All datasets generated and analyzed during this study are available upon request from the authors.
+Acknowledgments We acknowledge helpful discussions with O.
+Auslaender.
+This work was supported by the European Research Council (ERC-2024-SyG- 101167294 ; UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925, We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena.
+The Flatiron Institute is a division of the Simons Foundation.
+Support for hexagonal boron nitride crystal growth was provided by the Office of Naval Research, Award No.
+N00014-22-1-2582.
+K.K.
+acknowledges the support of the JSPJ KANENHI (21K18144), the JSPS Core-to-Core Program (JPJSCCA20240001) and the Alexander von Humboldt Foundation.
+The work at BNL was supported by the US Department of Energy, office of Basic Energy Sciences, contract no.
+DOE-SC0012704.
+D.M.
+and M.C.
+acknowledge support from the Gordon and Betty Moore Foundation’s EPiQS Initiative, Grant GBMF9069.
+D.M.K.
+and M.A.S.
+acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 531215165 (Research Unit ‘OPTIMAL’).
+M.A.S.
+was funded by the European Union (ERC, CAVMAT, project no.
+101124492).
+T.O.
+acknowledges support from the JSPS Overseas Research Fellowships.
+D.S.
+was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No.
+RS-2024-00333664).
+J.
+Y.
+and Q.L were supported by the Office of Basic Energy Sciences, Materials Sciences and Engineering Division, U.S.
+Department of Energy (DOE) under Contract No.
+DE-SC0012704.
+We acknowledge project/application support by the Max Planck Computing and Data Facility.
+Author contributions I.K., T.A.W., J.X., and G.P.
+collected and analyzed MFM data.
+I.K.
+and T.A.W.
+wrote the manuscript.
+S.Z.
+and S.S.Z.
+collected SNOM measurements.
+S.Z and J.Z.
+analyzed SNOM data.
+S.Z.
+simulated the HPhP dispersion for hBN on -ET.
+D.S.
+and B.S.Y.K.
+fabricated the heterostructures with input from C.R.D.
+G.P.
+performed SEM imaging of the MFM tips.
+J.Y.
+and Q.L.
+performed SQUID magnetometry measurements.
+T.O.
+provided simulation of the electric field in a z-dependent superfluid density scenario.
+J.H.E.
+grew the hBN.
+M.J., R.P., and S.W.
+provided discussion and interpretation of the data, as well as formulation of the idea.
+K.M.
+and K.K.
+grew the -ET crystals.
+G.G grew the BSCCO crystal.
+D.M.
+and M.C.
+grew the RuCl3 crystal.
+M.B.
+and A.C.
+provided discussion and interpretation of the data.
+D.M.K.
+provided guidance on fitting and participated in the discussion on mechanisms.
+D.S.
+performed the ab initio calculations under the supervision of A.R.
+D.S., M.H.M., and A.R.
+analyzed the ab initio results and developed a microscopic model for the quenching of superconductivity.
+M.A.S.
+provided theoretical input on the interpretation of results and contributed to the manuscript.
+A.N.P.
+advised on measurements and data interpretation.
+D.N.B.
+advised and supervised the project.
+All authors discussed the results and contributed to the final paper.
+Competing interests The authors declare no competing interests.
+Correspondence and requests for materials should be addressed to Itai Keren, Tatiana Webb, Shuai Zhang or Dmitri Basov.

### v4

@@
 Is it feasible to alter the ground state properties of a material by engineering its electromagnetic environment?
-Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [10, 11, 12, 13, 14, 15].
+Inspired by theoretical predictions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14], experimental realizations of such cavity-controlled properties without optical excitation are beginning to emerge [15, 16, 17, 18, 19, 20, 21].
 Here, we devised and implemented a novel platform to realize cavity-altered materials.
-Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [16, 17, 18, 19, 20, 21, 22].
+Single crystals of hyperbolic van der Waals (vdW) compounds provide a resonant electromagnetic environment with enhanced density of photonic states and prominent mode confinement [22, 23, 24, 25, 26].
 We interfaced hexagonal boron nitride (hBN) with the molecular superconductor -(BEDT-TTF)2Cu[N(CN)2]Br (-ET).
-The frequencies of infrared (IR) hyperbolic modes of hBN [23, 24] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [25, 26].
+The frequencies of infrared (IR) hyperbolic modes of hBN [27, 28] match the IR-active carbon-carbon stretching molecular resonance of -ET implicated in superconductivity [29].
 Nano-optical data supported by first-principles molecular Langevin dynamics simulations confirm the presence of resonant coupling between the hBN hyperbolic cavity modes and the carbon-carbon stretching mode in -ET.
 Meissner effect measurements via magnetic force microscopy demonstrate a strong suppression of superfluid density near the hBN/-ET interface.
-Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the superfluid suppression.
+Non-resonant control heterostructures, including RuCl3/-ET and hBN/, do not display the pronounced superfluid suppression.
 These observations suggest that hBN/-ET realizes a cavity-altered superconducting ground state.
-This work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties.
+Our work highlights the potential of dark cavities devoid of external photons for engineering electronic ground state properties of complex quantum materials.
 Some of the most exciting properties of solids arise from strong collective interactions among electrons, spins, and the crystal lattice.
@@
 At the vanguard of current interest and debate is the question whether the strong interaction of the quantum fluctuations of electromagnetic modes in photonic cavities or meta-structures with elementary excitations in solids can also prompt phase transitions and give rise to new quantum states of matter.
-A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9].
-Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 21, 7]; Amperean pairing directly mediated through dark-cavity photons [27, 28]; and the driven-cavity extension of the Eliashberg effect [29].
+A grand aspiration of cavity quantum materials research is to uncover fundamentally new routes for controlling properties of matter by judiciously tailoring the quantum electromagnetic environment [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
 Experiments with dark cavities revealed modified transport properties in the integer and fractional quantum Hall states of a 2D electron gas [15, 19, 21] as well as cavity-assisted thermal control of the metal-to-insulator transition in charge density wave systems [17].
-We focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings.
+Pioneering theoretical works on cavity control of superconductivity explored dark-cavity modification of phonon-mediated pairing through an increase in electron-phonon coupling via phonon polaritons [1, 10, 11], Amperean pairing directly mediated through dark-cavity photons [30, 31] as well as the driven-cavity extension of the Eliashberg effect [32].
+Here, we have chosen to focus on superconductivity as a purely electronic phase transition, devoid of lattice reconstructions and charge or spin orderings.
 We work with -(BEDT-TTF)2Cu[N(CN)2]Br (-ET), a widely studied layered organic salt that superconducts below the transition temperature =11.5 K [33, 34] (Supplementary Information Sections 1,2).
-We create an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN).
+We harness an electromagnetic environment structured by the dipole-active phonons of a thin van der Waals (vdW) hyperbolic material, hexagonal boron nitride (hBN).
 Hyperbolicity occurs when the permittivity has opposite signs along different axes [27, 28] and results in a highly enhanced photonic density of states [35, 36, 37].
-Notably, hyperbolic modes of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [35] and the two modes hybridize at the hBN/-ET interface.
-Effectively, the hBN slab acts as an electromagnetic cavity resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [25, 26].
-In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques that effectively analyze the electrodynamics of buried interfaces.
-Meissner force probe of resonant and non-resonant interfaces
-Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [37, 38, 39].
+Notably, hyperbolic modes (HMs) of hBN overlap with the frequency of the carbon-carbon (C=C) stretching mode of -ET at 1470 cm-1 [38] and the two modes hybridize at the hBN/-ET interface.
+Effectively, the hBN slab acts as an electromagnetic cavity (Supplementary Information Section 3) resonantly tuned to the C=C stretching mode of -ET that is implicated in superconductivity [29].
+In this work, we examine how superconductivity is altered at the hBN/-ET interface, using advanced scanning probe-based techniques adept at the analysis of electrodynamics in buried interfaces.
+Meissner effect at resonant and non-resonant interfaces
+Magnetic force microscopy (MFM) probes the local superfluid density via the repulsive Meissner force experienced by a magnetized tip at the end of a cantilever oscillating above the surface of a superconductor [41, 42, 43, 44].
 The force results from the supercurrent screening the stray magnetic field of the tip.
@@
 The magnitude of grows with reduced tip-sample surface separation.
-The signal collected above hBN (blue) shows a weakened Meissner force, highlighted by the gray shading.
-As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of -ET.
+The signal collected above hBN (blue) shows a weakened , highlighted by the gray shading.
+As a control experiment, we placed a 55-nm-thick RuCl3 microcrystal on top of the same -ET sample.
 RuCl3 is an insulator with a static permittivity similar to that of hBN [46, 40].
@@
 The stark contrast between the RuCl3/-ET and hBN/-ET interfaces indicates that the suppression of superfluid density under hBN is likely prompted by resonant coupling across the hBN/-ET interface.
-We remark that the RuCl3 control experiment rules out a prominent role for work function disparity [42] in the suppression of the superfluid density in Fig.
-1 and similarly suggests that strain is unlikely to be significant.
+We remark that the RuCl3 control experiment rules out a prominent role of charge transfer due to work function disparity [47] in the suppression of the superfluid density in Fig.
+1 (Supplementary Information Section 14) and similarly suggests that strain is unlikely to be significant.
 We perform a second control measurement on a heterostructure of 8-nm hBN residing on a 28.5-nm-thick microcrystal of (BSCCO) with K.
-Crystals of BSCCO host phonons below 650 cm-1 [43], far below the hyperbolic modes in hBN (Supplementary Information Section 5).
+Crystals of BSCCO host phonons below 650 cm-1 [48], far below the HMs in hBN.
 MFM data plotted in Fig.
-1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray).
+1b (inset) for hBN/BSCCO (blue) are indistinguishable from the data for bare BSCCO (dark gray - Supplementary Information Section 6).
 Thus, the significant suppression of the superfluid density is specific to the resonant nature of the hBN/-ET interface.
-Electrodynamics of the hBN/-ET interface
-Next, we investigate resonant mode coupling across the hBN/-ET interface.
-The hBN crystal hosts hyperbolic phonon polaritons (HPhPs) that exhibit light-like dispersion near the transverse optical (TO) frequency and phonon-like dispersion near the longitudinal optical (LO) frequency [16].
-Here, TO and LO refer specifically to the in-plane modes.
-We examine the HPhP dispersion upon interfacing with -ET through the calculated imaginary part of the p-polarized Fresnel reflection , as shown in Fig.
-2a (Supplementary Information Section 6).
-For free-standing hBN, the HPhP dispersion has a series of branches between the TO and LO phonon frequencies, each with a continuous dispersion [23].
-By contrast, for an hBN/-ET heterostructure, the HPhP branches are interrupted near the C=C stretching frequency.
-Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully opened due to broadening caused by dielectric loss associated with the in-plane metallic response of -ET [45, 46].
-To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [47].
-Specifically, we evaluated the hBN HPhP dispersion as these modes couple to the C=C stretching mode of -ET.
-We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser.
-The strong electric field with high momentum achieved near the tip apex launches polariton waves.
-We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET.
-Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig.
-2b for cm-1.
-Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig.
-2c).
-In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [48].
-However, in our data on -ET-supported hBN, the fringe evolution exhibits a kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig.
-2c.
-These nano-IR data attest to mode coupling across the hBN--ET interface.
-To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HPhP and the C=C stretching mode.
-First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the HPhP’s electric field with the C=C stretching mode.
-The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig.
-2d and Supplementary Information Section 2).
-It is important to note that the HPhP field has an out-of-plane component that can couple to the C=C mode (Fig.
-2d).
-The combined contributions of many HPhP branches outspreading to high momenta make this field extend far beyond the hBN surface.
-We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1).
-Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 7) shows that the out-of-plane electric field induced by the zero-point HPhP fluctuations reduces the C=C amplitude and splits the peak in Fig.
-2e.
-We therefore conclude that vacuum fluctuations—specifically, zero-point phonon motion—mediate the coupling between hyperbolic phonon polaritons and molecular vibrations without the need for external photons, thus enabling interaction in the dark.
 Suppression of superfluid density at the hBN/-ET interface
@@
 1a) with suppressed superfluid density by studying the temperature dependence.
-Over bare -ET, the large repulsive is rooted in the Meissner effect (Supplementary Information Section 4), it decreases with increasing temperature (Fig.
-3a) and vanishes near .
-Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces by performing constant-height scans with the tip hovering above the -ET surface in a 1616-m2 region that includes both hBN and RuCl3 (Fig.
-3b-d).
-To facilitate comparison between temperatures, we focus on the differential signal , where the latter term stands for the Meissner signal from bare -ET.
-At 2 K (panel b), pN/m over hBN, indicating a weaker Meissner force.
-As with the curves in Fig.
-1(a,b), the effect over RuCl3 is negligible.
-Warming to 12 K, just above , the image contrast disappears, giving everywhere.
-The vanishing near confirms that the weakened observed at 2 K is associated with a reduction in the superfluid density.
-Further evaluation and additional control measurements are described in Supplementary Information Section 8.
+Over bare -ET, the large repulsive arises from the Meissner effect (Supplementary Information Section 5), decreases with increasing temperature (Fig.
+2a), and vanishes near .
+Having confirmed that Meissner repulsion is the main contributor to over bare -ET, we examine the effects at the heterostructure interfaces.
+We performed constant-height scans with the tip hovering above the -ET surface in a 1616 m2 region that includes both hBN and RuCl3 (Fig.
+2b,c).
+To facilitate comparison between data collected at different temperatures, we focus on the differential signal , where the latter term stands for over bare -ET.
+At 2 K (Fig.
+2b), pN/m over hBN, indicating suppressed superfluid density.
+Over RuCl3, is visibly much smaller.
+More quantitatively, in Fig.
+2d we display the histogram corresponding to the image of the signals.
+We witness distinct distributions representing bare -ET, RuCl3/-ET, and hBN/-ET.
+The hBN/-ET peak is prominently separated from the other two, indicating the strong effect of hBN on the superfluid density of -ET.
+Supplementary Information Sections 7 and 8 discuss additional effects, including possible scenarios for the minor difference between the MFM data registered at the RuCl3/-ET interface and the bare -ET.
+Upon warming to 12 K, just above , the image contrast disappears, giving everywhere (Fig 2c).
+The vanishing near confirms that the weakened signal observed at 2 K is associated with a reduction in the superfluid density.
 We observe suppression of superfluid density under hBN in multiple hBN/-ET heterostructures.
@@
 The results are summarized in Fig.
-4.
-To model the Meissner force analytically [37], we make the simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, .
+3.
+To model analytically [41], we make the grossly simplifying assumption that the suppression of the superfluid density is uniform in depth and evaluate the effective superfluid density under hBN, .
 In Fig.
-4, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses.
+3, we compile MFM data for several devices with hBN cavities and RuCl3 dielectric layers of varied thicknesses.
 We present data in the form of , where is the superfluid density in bare -ET.
@@
 Accurate conversion between and superfluid density would require knowledge of the depth profile (Supplementary Information Section 5).
-While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar Meissner forces.
-However, the magnitude of the observed Meissner force suppression implies that -ET is impacted by the cavity to a significant depth.
-The impacted depth appears to be of the same order of magnitude as the polaritonic wavelength (Supplementary Information Section 10).
+While we expect the suppression of superfluid density to be strongest in the immediate vicinity of hBN and to decay into the -ET bulk, MFM measurements alone cannot differentiate among the possible depth profiles that all lead to similar .
+Nevertheless, the magnitude of the observed suppression of implies that superconductivity in -ET is impacted by the cavity to a significant depth (Supplementary Information Section 9).
+Electrodynamics of the hBN/-ET interface
+Next, we investigate resonant mode coupling across the hBN/-ET interface.
+The hBN crystal hosts HMs that exhibit light-like dispersion near the transverse optical (TO) frequency and phonon-like dispersion near the longitudinal optical (LO) frequency [22].
+Here, TO and LO refer specifically to the in-plane modes.
+In practice, the HMs manifest as hyperbolic phonon-polaritons (HPhPs) that can be visualized in nano-optical experiments as originally done in Ref.
+[27].
+We examine the HPhP dispersion of hBN upon interfacing with -ET through the calculated imaginary part of the p-polarized Fresnel reflection , as shown in Fig.
+4a (Supplementary Information Section 10).
+For free-standing hBN, the HPhP dispersion displays a series of branches between the TO and LO phonon frequencies, each with a continuous dispersion [27].
+By contrast, for an hBN/-ET heterostructure, the HPhP branches are interrupted near the C=C stretching frequency (light blue box in Fig.
+4a).
+Avoided crossings are visible, yet the gaps in the HPhP dispersion are not fully developed due to broadening caused by dielectric loss associated with the in-plane response of -ET [50, 51] (Supplementary Information Section 10).
+To experimentally examine the mode coupling at the hBN/-ET interface, we used scattering-type scanning near-field optical microscopy (s-SNOM) [52].
+Specifically, we evaluated the HPhP dispersion of hBN as its modes couple to the C=C stretching mode of -ET.
+We illuminated the metallized tip of an atomic force microscope with a frequency-tunable IR laser.
+The strong electric field with high momentum achieved near the tip apex launches polaritonic waves.
+We performed nano-IR measurements at K on a region of an hBN/-ET heterostructure with 32-nm hBN on 100-nm -ET.
+Fringes arising from the interference of propagating HPhPs were observed in the nano-IR images: representative results are shown in Fig.
+4b for cm-1.
+Line profiles from s-SNOM collected with varying incident light frequency show that the HPhP wavelength changes with photon energy (Fig.
+4c).
+In free-standing hBN, the real-space fringes arising from HPhP interference evolve smoothly with the frequency [53].
+However, in our data on -ET-supported hBN, the fringe evolution exhibits an evident kink near the C=C frequency, indicated by the solid line tracking the fifth peak in Fig.
+4c.
+These nano-IR data attest to mode coupling across the hBN--ET interface
+To gain insight into the experimentally observed mode coupling, we take a closer look at the interplay between the zero-point fluctuations of the hBN HMs and the C=C stretching mode.
+First-principles molecular Langevin dynamics simulations indicate that the coupling arises from the interaction of the transverse component of the electric field of the HMs with the C=C stretching mode.
+The C=C stretching mode exhibits a dipole moment that is directed primarily out-of-plane (Fig.
+4d and Supplementary Information Section 2).
+It is important to note that the electric field of the HM has an out-of-plane component that can couple to the C=C mode (Fig.
+4d).
+The combined contributions of many HM branches reaching high momenta make this field extend far beyond the hBN surface (Supplementary Information Section 3).
+We therefore calculate the C=C amplitude in the presence of an oscillating out-of-plane electric field with a frequency spectrum spanning the C=C frequency (fundamental frequency cm-1).
+Langevin molecular dynamics analysis carried out at K (Supplementary Information Section 11) shows that the out-of-plane electric field induced by the zero-point HM fluctuations reduces the C=C amplitude and splits the peak in Fig.
+4e.
+We therefore conclude that zero-point HM motion mediates the coupling between HMs and molecular vibrations without the need for external photons.
 Hyperbolic cavity electrodynamics
 The data in Figs.
-1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium property of superconductivity.
-These experiments establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering.
-Our specific implementation of the resonant structure reveals two notable features of hyperbolic cavity control.
-First, hyperbolic modes entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields.
-The out-of-plane component is crucial for coupling to the out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation.
-Second, the multiple branches of hyperbolic modes span a broad momentum range, enabling continuous interaction with the weakly dispersing C=C resonance across the entire phase space, as shown in Fig.
-2a.
-Future experiments will be needed to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [49].
-Our data and modeling suggest a plausible pathway for cavity enhancement of superconducting properties.
-Indeed, our simulations indicate that the hBN cavity subdues the C=C stretching mode fluctuations prompting the stark suppression of superfluid density.
-By the same token, boosting fluctuations through mode softening in the cavity [1] may increase the superfluid density and possibly in -ET.
-Furthermore, the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulations of material properties is rather general.
-We remind the reader that vdW crystals host hyperbolic modes of phononic, plasmonic and exitonic origin [50] spanning the range from THz to visible, thus offering limitless options for hyper-cavity engineering.
+1-4 demonstrate that cavity electrodynamics plays a pivotal role in suppressing the superfluid density – a hallmark equilibrium attribute of superconductivity.
+These experiments also establish a precedent for successfully altering a material’s ground state thermodynamic property through cavity engineering.
+Our specific implementation of the resonant structure exploits two notable features of hyperbolic cavity control.
+First, HMs entail both in-plane and out-of-plane components of the fluctuating electromagnetic fields.
+The out-of-plane component is crucial for coupling to the primarily out-of-plane dipole moment of the C=C stretching mode in the -ET superconductor under investigation.
+Second, the multiple branches of the HMs enable interaction with the weakly dispersing C=C resonance at multiple intersections over a broad range of momenta, as shown in Fig.
+4a.
+Future experiments need to evaluate the spatial extent of the mode coupling away from the hyperbolic interface; our earlier data for a similar resonant hBN/MoO3 heterostructure suggest that the relevant length scale may exceed nm [54].
+Our data and modeling suggest a plausible pathway for cavity enhancement of superconductivity.
+As an illustrative example, we consider the scenario of an optical phonon in a superconductor coupled to a hBN cavity within the framework of Bardeen, Cooper and Schrieffer (BCS) theory (Supplementary Information Section 13).
+Electron-phonon coupling, quadratic in the phonon coordinate [1, 55], results in a suppression of the superfluid density, while linear coupling results in enhancement.
+This analysis suggests that conventional superconductors may exhibit increased superfluid density through interactions with HMs.
+Here we pause to caution the reader: -ET studied in our work is an unconventional superconductor [56], and BCS-derived inferences can only guide our intuition.
+A rigorous description of the physics involved in cavity-coupled -ET awaits further investigation.
+We conclude by highlighting that the concept of employing readily reconfigurable vdW hyperbolic cavities for resonant manipulation of material properties is rather general.
+HMs of phononic, plasmonic and excitonic origin are hosted in many vdW materials including insulators, semiconductors as well as a variety of emerging materials [26].
+These modes span the range from THz to visible [57], thus offering limitless options for hyper-cavity engineering.
 - [1] Sentef, M.
@@
 - [2] Ruggenthaler, M., Tancogne-Dejean, N., Flick, J., Appel, H. & Rubio, A. From a quantum-electrodynamical light–matter description to novel spectroscopies. Nat. Rev. Chem. 2, 1–16 (2018).
+- [3] Flick, J., Rivera, N. & Narang, P. Strong light-matter coupling in quantum chemistry and quantum photonics. Nanophotonics 7, 1479–1501 (2018).
 - [4] Garcia-Vidal, F. J., Ciuti, C. & Ebbesen, T. W. Manipulating matter by strong coupling to vacuum fields. Science 373, eabd0336 (2021).
@@
 - [6] Schlawin, F., Kennes, D. M. & Sentef, M. A. Cavity quantum materials. Appl. Phys. Rev. 9 (2022).
+- [7] Bloch, J., Cavalleri, A., Galitski, V., Hafezi, M. & Rubio, A. Strongly correlated electron–photon systems. Nature 606, 41–48 (2022).
+- [8] Ebbesen, T. W., Rubio, A. & Scholes, G. D. Introduction: Polaritonic Chemistry. Chem. Rev. 123, 12037–12038 (2023).
 - [9] Viñas Boström, E., Sriram, A., Claassen, M. & Rubio, A. Controlling the magnetic state of the proximate quantum spin liquid -RuCl3 with an optical cavity. npj Comput. Mater. 9, 202 (2023).
+- [10] Hübener, H., Boström, E. V., Claassen, M., Latini, S. & Rubio, A. Quantum materials engineering by structured cavity vacuum fluctuations. Mater. Quantum. Technol. 4, 023002 (2024).
 - [11] Lu, I.-T. et al. Cavity-enhanced superconductivity in MgB2 from first-principles quantum electrodynamics (QEDFT). PNAS 121, e2415061121 (2024).
-- [8] Flick, J., Rivera, N. & Narang, P. Strong light-matter coupling in quantum chemistry and quantum photonics. Nanophotonics 7, 1479–1501 (2018).
 - [12] Lu, I.-T. et al. Cavity engineering of solid-state materials without external driving. Adv. Opt. Photon. 17, 441–525 (2025).
+- [13] Fassioli, F., Faist, J., Eckstein, M. & Fausti, D. Controlling radiative heat flow through cavity electrodynamics. Phys. Rev. B 111, 165425 (2025).
+- [14] Riolo, R. et al. Tuning Fermi liquids with polaritonic cavities. PNAS 122, e2407995122 (2025).
 - [15] Paravicini-Bagliani, G. et al. Magneto-transport controlled by landau polariton states. Nat. Phys. 15, 186–190 (2019).
+- [16] Appugliese, F. et al. Breakdown of topological protection by cavity vacuum fields in the integer quantum Hall effect. Science 375, 1030–1034 (2022).
 - [17] Jarc, G. et al. Cavity-mediated thermal control of metal-to-insulator transition in 1T-TaS2. Nature 622, 487–492 (2023).
@@
 - [24] Basov, D. N., Asenjo-Garcia, A., Schuck, P. J., Zhu, X. & Rubio, A. Polariton panorama. Nanophotonics 10, 549–577 (2020).
-- [19] Bloch, J., Cavalleri, A., Galitski, V., Hafezi, M. & Rubio, A. Strongly correlated electron–photon systems. Nature 606, 41–48 (2022).
-- [20] Ebbesen, T. W., Rubio, A. & Scholes, G. D. Introduction: Polaritonic Chemistry. Chem. Rev. 123, 12037–12038 (2023).
-- [21] Hübener, H., Boström, E. V., Claassen, M., Latini, S. & Rubio, A. Quantum materials engineering by structured cavity vacuum fluctuations. Mater. Quantum. Technol. 4, 023002 (2024).
 - [25] Herzig Sheinfux, H. et al. High-quality nanocavities through multimodal confinement of hyperbolic polaritons in hexagonal boron nitride. Nat. Mater. 23, 499–505 (2024).
+- [26] Basov, D. et al. Polaritonic quantum matter. Nanophotonics (2025).
 - [27] Dai, S. et al. Tunable phonon polaritons in atomically thin van der waals crystals of boron nitride. Science 343, 1125–1129 (2014).
@@
 - [29] Buzzi, M. et al. Photomolecular High-Temperature Superconductivity. Phys. Rev. X. 10, 31028 (2020).
-- [26] Buzzi, M. et al. Phase Diagram for Light-Induced Superconductivity in . Phys. Rev. Lett. 127, 197002 (2021).
 - [30] Schlawin, F., Cavalleri, A. & Jaksch, D. Cavity-mediated electron-photon superconductivity. Phys. Rev. Lett. 122, 133602 (2019).
@@
 - [34] Dressel, M. et al. Electrodynamics of the organic superconductorsinebreak -(BEDT-TTF)2Cu(NCS)2 and -(BEDT-TTF)2Cu[N(CN)2]Br. Phys. Rev. B 50, 13603 (1994).
-- [32] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photonics 7, 948–957 (2013).
+- [35] Poddubny, A., Iorsh, I., Belov, P. & Kivshar, Y. Hyperbolic metamaterials. Nat. Photon. 7, 948–957 (2013).
 - [36] Sun, J., Litchinitser, N. M. & Zhou, J. Indefinite by Nature: From Ultraviolet to Terahertz. ACS Photonics 1, 293–303 (2014).
@@
 - [38] McGuire, J. J. et al. Incoherent interplane conductivity of . Phys. Rev. B 64, 094503 (2001).
+- [39] Giles, A. et al. Ultralow-loss polaritons in isotopically pure boron nitride. Nat. Mater. 17, 134–139 (2018).
 - [40] Reschke, S. et al. Electronic and phonon excitations in -RuCl3. Phys. Rev. B 96, 165120 (2017).
@@
 - [43] Lamhot, Y. et al. Local characterization of superconductivity in . Phys. Rev. B 91, 060504 (2015).
+- [44] Grebenchuk, S. Y. et al. Crossover from ferromagnetic superconductor to superconducting ferromagnet in P-doped EuFe2(As1-xPx)2. Phys. Rev. B 102, 144501 (2020).
 - [45] Hartmann, U. Magnetic Force Microscopy. Annu. Rev. Mater. Sci. 29, 53–87 (1999).
@@
 - [51] Faltermeier, D. et al. Bandwidth-controlled mott transition in : Optical studies of localized charge excitations. Phys. Rev. B 76, 165113 (2007).
-- [47] Chen, X. et al. Modern scattering-type scanning near-field optical microscopy for advanced material research. Adv. Mater. 31, 1804774 (2019).
-- [48] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photonics 15, 197–202 (2021).
+- [52] Hillenbrand, R., Abate, Y., Liu, M., Chen, X. & Basov, D. N. Visible-to-THz near-field nanoscopy. Nat. Rev. Mater. 10, 285–310 (2025).
+- [53] Bylinkin, A. et al. Real-space observation of vibrational strong coupling between propagating phonon polaritons and organic molecules. Nat. Photon. 15, 197–202 (2021).
 - [54] Sternbach, A. J. et al. Negative refraction in hyperbolic hetero-bicrystals. Science 379, 555–557 (2023).
-- [50] Basov, D. et al. Polaritonic quantum matter. Nanophotonics (2025).
+- [55] Kennes, D. M., Wilner, E. Y., Reichman, D. R. & Millis, A. J. Transient superconductivity from electronic squeezing of optically pumped phonons. Nat. Phys. 13, 479–483 (2017).
+- [56] Buzzi, M. et al. Phase Diagram for Light-Induced Superconductivity in . Phys. Rev. Lett. 127, 197002 (2021).
+- [57] Ruta, F. L. et al. Good plasmons in a bad metal. Science 387, 786–791 (2025).
 Methods
 Device fabrication
-Heterostructures were fabricated by placing vdW materials, on the order of 10 nm to 120 nm thick, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2).
-Measurements were performed on five devices (Supplementary Information Section 3).
+The heterostructures were fabricated by placing vdW materials of thicknesses on the order of 10 nm to 120 nm, on large flat surfaces of -ET single crystals (ac crystallographic plane, Supplementary Information Section 2).
+Measurements were performed on six devices (Supplementary Information Section 4).
 Data in Fig.
@@
 4 from Device 2, in Fig.
-3(b-d) from device 1, and in Fig.
-3a from device 5.
-Fig.
-4 is a compilation of results from devices 1, 3, and 4.
-The -ET crystal surfaces were either as-grown (devices 1 and 3), exposed by cleaving bulk crystals (devices 4 and 5), or exposed by exfoliating single crystals (device 2).
+2(b-d) from Device 1, and in Fig.
+2a from Device 6.
+The data in Fig.
+3 represent a compilation of results from Devices 1, 3, 4, and 5.
+The -ET crystal surfaces were either as-grown (Devices 1, 4, 5), exposed by cleaving bulk crystals (Devices 3 and 6), or exposed by exfoliating single crystals (Device 2).
 For bulk samples, the -ET single crystal was fixed onto a silicon substrate using either epoxy or silver paint.
-For devices 4 and 5, the crystal was then cleaved to expose new surfaces.
+For Devices 3 and 6, the crystal was then cleaved to expose new surfaces.
 Large atomically flat surfaces were then selected by optical and atomic force microscopy.
@@
 Cleaved surfaces showed larger variations in surface roughness.
-Both hBN and RuCl3 flakes were exfoliated using a standard scotch-tape technique onto a polydimethylsiloxane (PDMS) stamp and identified via optical and atomic force microscopy.
+For each device, MFM measurements spanned time periods ranging from weeks to months, and in general included multiple thermal cycles.
+For all devices, no sign of degradation was found.
+Both hBN and RuCl3 microcrystals were exfoliated using a standard scotch-tape technique onto a polydimethylsiloxane (PDMS) stamp and identified via optical and atomic force microscopy.
 The stacking process was performed using a commercial transfer stage with temperature control.
 The PDMS stamp was gently pressed onto the -ET surface and then slowly retracted.
-The hBN flakes were released at approximately 60°C, while the RuCl3 flakes were released at around 110°C.
+The hBN microcrystals were released at approximately 60°C, while the RuCl3 microcrystals in Devices 1 and 5 were released at around 110°C and 130°C, respectively.
 Isotopically pure hBN (10BN) was used to achieve a narrow linewidth for the TO phonon.
@@
 BSCCO was first exfoliated on SiO2 substrates.
-hBN flakes were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
+hBN microcrystals were then exfoliated on a PDMS stamp and transferred onto BSCCO at room temperature.
 s-SNOM
@@
 The average curves were then smoothed with a uniform filter with width 30 to 50 pixels to reduce the appearance of noise.
-For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the Meissner force is negligible.
+For the -ET heterostructures, a linear background was subtracted by fitting the data far from the sample surface where the is negligible.
 For the BSCCO sample, a background measured on the silicon substrate was subtracted (Supplementary Information Section 6).
@@
 Auslaender.
+The authors acknowledge the use of facilities and instrumentation supported by NSF through the Columbia University, Columbia Nano Initiative, and the Materials Research Science and Engineering Center DMR-2011738.
+D.N.B is a Moore Investigator in Quantum Materials EPIQS GBMF9455.
+Research at Columbia on electrodynamics of van der Waals interfaces is supported by DOE-BES DE-SC0018426.
+Fabrication of vdW heterostructures at Columbia and the development of nano-optical methods is supported as part of Programmable Quantum Materials, an Energy Frontier Research Center funded by the U.S.
+Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under award DE-SC0019443.
+Advances in MFM imaging at Columbia are supported by ARO award W911NF2510062.
 This work was supported by the European Research Council (ERC-2024-SyG- 101167294 ; UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925, We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena.
@@
 T.O.
-provided simulation of the electric field in a z-dependent superfluid density scenario.
+provided simulations of the electric field in a z-dependent superfluid density scenario, and in an oscillating dipole above hBN scenario.
 J.H.E.
@@
 provided guidance on fitting and participated in the discussion on mechanisms.
+M.H.M.
+computed the quantum electric field fluctuations from a thin film of hBN.
 D.S.
@@
 analyzed the ab initio results and developed a microscopic model for the quenching of superconductivity.
+E.V.B.
+developed the analytical model for BCS-HM coupling.
 M.A.S.